### v1

# ReaCritic: Large Reasoning Transformer-based DRL Critic-model Scaling For Heterogeneous Networks

Abstract
Heterogeneous Networks (HetNets) pose critical challenges for intelligent management due to the diverse user requirements and time-varying wireless conditions. These factors introduce significant decision complexity, which limits the adaptability of existing Deep Reinforcement Learning (DRL) methods. In many DRL algorithms, especially those involving value-based or actor-critic structures, the critic component plays a key role in guiding policy learning by estimating value functions. However, conventional critic models often use shallow architectures that map observations directly to scalar estimates, limiting their ability to handle multi-task complexity. In contrast, recent progress in inference-time scaling of Large Language Models (LLMs) has shown that generating intermediate reasoning steps can significantly improve decision quality. Motivated by this, we propose ReaCritic, a large reasoning transformer-based critic-model scaling scheme that brings reasoning ability into DRL. ReaCritic performs horizontal reasoning over parallel state-action inputs and vertical reasoning through deep transformer stacks. It is compatible with a broad range of value-based and actor-critic DRL algorithms and enhances generalization in dynamic wireless environments. Extensive experiments demonstrate that ReaCritic improves convergence speed and final performance across various HetNet settings and standard OpenAI Gym control tasks.
Index Terms:
Transformer, heterogeneous networks, optimization, reasoning, and large language models
I Introduction
The rapid growth of intelligent wireless services such as the Internet of Things (IoT), Mobile Edge Computing (MEC), and smart city applications is driving the evolution of future wireless networks toward large-scale Heterogeneous Networks (HetNets) [1]. These HetNets involve a large number of mobile users and dynamic environments with heterogeneous communication and computing resources. As a result, HetNets must handle massive data traffic, ensure low-latency responses, and maintain energy efficiency under diverse service demands, which introduces fundamental challenges for resource management and decision-making [2]. In particular, the coexistence of heterogeneous user types, diverse Quality-of-Service (QoS) requirements, and time-varying wireless conditions leads to high-dimensional state spaces, complex user-resource interactions, and non-stationary dynamics. These properties hinder the scalability and generalization of traditional optimization and supervised learning-based methods [3].
Figure 1: Comparison of DRL critic designs for HetNets. Conventional MLP-based critics (Part A) struggle with generalization in high-dimensional spaces. LLM-aided critics (Part B) improve expressiveness but face domain mismatch and integration challenges. Our proposed ReaCritic (Part C) introduces transformer-based bidirectional reasoning to achieve more accurate value estimation and scalable DRL training.
Faced with these challenges, Deep Reinforcement Learning (DRL) has gained attention as a data-driven approach for autonomous network management in complex wireless environments. By interacting with the environment and optimizing long-term rewards, DRL enables agents to adapt to dynamic network conditions without relying on predefined models [4]. Its success in domains such as intelligent control, robotics, and wireless resource management demonstrates its potential for handling large-scale, uncertain systems [5, 6, 7]. However, when applied to large-scale HetNets, DRL still faces key limitations not only due to environmental complexity but also due to internal architectural bottlenecks. Although DRL algorithms are capable of learning adaptive behaviors through trial-and-error interaction, many existing designs rely on simple Multi-Layer Perceptron (MLP)-based model structures that treat input states as flat feature vectors and apply direct approximations for policy or value functions. This formulation fundamentally lacks the capacity to capture latent dependencies among state variables or to represent tradeoffs inherent in multi-objective decision-making. In heterogeneous wireless environments, where metrics such as latency, throughput, and energy efficiency are often tightly coupled and evolve asynchronously, such limitations can lead to unstable training, slow convergence, and poor performance in dynamic or unseen scenarios. Taking actor-critic-based DRL algorithms as an example, as shown in Fig. 1, the critic network estimates Q-values based on environmental states and the actions selected by the actor network. However, due to the high dimensionality and structural complexity of both the HetNet states and the action space, this mapping becomes difficult and inaccurate, limiting the effectiveness of actor updates and ultimately degrading overall DRL performance.
Therefore, it is essential to enhance the scalability of DRL frameworks to better accommodate high-dimensional state spaces and dynamic network behaviors. One common strategy is to increase model depth or width, with the goal of improving the DRL’s capacity to interpret complex input features. However, scaling up the MLP parameter size alone often leads to overfitting, unstable convergence, and limited robustness in non-stationary environments. A more effective direction is to improve the DRL model through architectural adjustments that enable it to process structured observations and capture latent dependencies within the environment, beyond what can be achieved through conventional model size expansion. In this context, the architectural principles of Large Language Models (LLMs) offer a compelling source of inspiration. Recent advances in reasoning-capable LLMs have demonstrated strong capabilities in pattern abstraction and generalization across a wide range of tasks [8]. These LLMs are particularly effective at modeling intricate relationships and long-term dependencies in structured or semi-structured data, opening new opportunities for improving representation learning and decision-making in complex environments.
Motivated by the capabilities of LLMs, recent research has explored LLM-aided DRL frameworks, where well-pretrained LLMs are employed to assist in policy generation, reward shaping, or action refinement [9, 10, 11, 12]. While these approaches have shown promise in certain domains, directly incorporating LLMs into the DRL training loop introduces several critical limitations. First, LLMs are primarily designed for natural language understanding and generation, which does not naturally align with the optimization-driven, reward-sensitive objectives of DRL. Second, LLMs exhibit non-deterministic behavior that is often difficult to control, resulting in unstable learning dynamics and inconsistent policy updates [13]. Third, integrating large-scale LLMs incurs substantial computational overhead and lacks theoretical support regarding convergence and stability in control settings [14]. These limitations indicate that enhancing the representation and reasoning capacity of DRL through architectural design remains a key challenge:
  • Q1 (How): How can the reasoning capability of LLMs be effectively adapted to DRL settings, if not directly incorporating pre-trained LLMs as external modules?
  • Q2 (Where): Which DRL component is most suitable for integrating LLM-inspired reasoning and scaling to support complex, dynamic environments?
To address the above questions, it is necessary to first reflect on the source of LLMs’ reasoning capabilities. Fundamentally, these capabilities do not arise from language data alone, but from the architectural properties of Transformer networks. Through self-attention and hierarchical composition, Transformers can capture relational dependencies and long-range context, forming the backbone of scalable reasoning in LLMs. This insight suggests that one can extract and adapt the structural mechanisms underlying LLMs’ reasoning processes. Recent works have explored Transformer-based designs to improve value estimation, policy modeling, and trajectory learning in complex environments [15, 16]. For instance, the authors in [15] proposed a lightweight Transformer model for autonomous driving that combines temporal semantic extraction and risk modeling to generate safer driving strategies. Similarly, the authors in [16] developed SACCT, a DRL framework for adaptive live streaming, where a communication Transformer captures multi-scale dependencies to jointly optimize uplink scheduling and edge transcoding. However, existing approaches primarily focus on task-specific modeling enhancements and do not explicitly introduce structured reasoning processes that support scalable, general-purpose decision-making across dynamic network environments.
To address Q1 (How), we adopt a Transformer-based reasoning architecture as a more stable, scalable, and task-adaptive alternative to conventional MLP-based designs. By leveraging self-attention mechanisms and stacked abstraction layers, the Transformer enables structured reasoning over state-action representations. This allows for more accurate value estimation under multi-objective and dynamic conditions, while avoiding the overfitting and instability often associated with naive MLP scaling. The hierarchical and modular design of the Transformer also facilitates its integration into DRL training pipelines, where robustness and adaptability are essential. To address Q2 (Where), we embed this reasoning capability into the critic module of the DRL framework. As the critic is responsible for value estimation and provides gradient signals to guide policy updates, it is the natural component for incorporating structured inference. This choice preserves the original policy structure while substantially improving the agent’s ability to model long-term dependencies, reason over complex state-action trajectories, and adapt to heterogeneous and time-varying network environments. Our proposed design, referred to as ReaCritic, is illustrated in Part C of Fig. 1.
The main contributions of this paper are as follows:
  • We formulate a DRL-based resource management problem for large-scale HetNets, capturing the joint effects of heterogeneous user types, time-varying wireless conditions, and asynchronous service demands. This formulation reflects key challenges such as non-stationarity and high dimensionality, motivating the need for scalable DRL frameworks with enhanced reasoning capabilities.
  • We propose ReaCritic, a large reasoning transformer-based critic-model scaling architecture for DRL, inspired by inference-time scaling in LLMs. ReaCritic enables horizontal reasoning across parallel state-action token sequences and vertical reasoning through stacked transformer blocks, supporting expressive and adaptable Q value estimation in complex environments.
  • We conduct extensive experiments in a large-scale multi-user HetNet environment to evaluate ReaCritic. Compared with conventional MLP-based critics and standard DRL baselines, ReaCritic achieves faster convergence, improved policy stability, and stronger adaptability under dynamic and heterogeneous network conditions.
The remainder of this paper is organized as follows: Section II discusses related work in HetNets, DRL for networks, and transformer for wireless networks. Section III presents the wireless network system model. Section III-B gives the problem formulation of the optimization problem in the considered HetNet. Section IV details the design of ReaCritic and the reasoning transformer-based critic-model scaling. Section V presents comprehensive experimental results and analysis. Finally, Section VI concludes the paper with a summary of our key findings. A list of mathematical symbols frequently used in this paper is shown in Table I.
TABLE I: Mathematical Notations
Notation Description
$x_{m}$ Position of the $m_{\rm th}$ user
$d_{m}$ Computation demand (in CPU cycles)
$f_{m}$ Local CPU frequency of the $m_{\rm th}$ user
$p_{\mathrm{u},m},p_{\mathrm{d},m}$ Normalized uplink/downlink power allocation factors
$b_{\mathrm{u},m},b_{\mathrm{d},m}$ Normalized uplink/downlink bandwidth allocation factors
$c_{m}$ Normalized computation resource allocation factor
$q_{\mathrm{u},m},q_{\mathrm{d},m}$ Uplink/downlink Rician fading gain
$\sigma_{\mathrm{u}}^{2},\sigma_{\mathrm{d}}^{2}$ Uplink/downlink noise power
$P_{\mathrm{u},m}$ Maximum uplink transmit power of the $m_{\rm th}$ user
$P_{\mathrm{d},\max}$ Maximum downlink transmit power of the BS
$B_{\mathrm{u},m},B_{\mathrm{d},m}$ Maximum uplink/downlink bandwidth
$\zeta_{m}$ Large-scale path loss of the $m_{\rm th}$ user
$C_{\max}$ Total available computation capacity
$\rho_{m}$ Local compute efficiency of the $m_{\rm th}$ user
$\tau_{m}$ User type index
$Q_{m}$ Utility function of the $m_{\rm th}$ user
$\alpha_{m}^{(i)}$ Weight for the $i_{\rm th}$ objective
$\mathcal{R}_{\mathrm{u},m},\mathcal{R}_{\mathrm{d},m}$ Uplink/downlink data rate
$\eta_{\mathrm{u},m},\eta_{\mathrm{d},m}$ Uplink/downlink energy efficiency
$L_{m}$ Service latency
$l_{\mathrm{th},m}$ Latency threshold for the $m_{\rm th}$ user
$B$ Batch size of DRL
$d_{\rm h}$ The hidden dimension size
$H$ The number of horizontal reasoning steps
$V$ The number of vertical reasoning steps, i.e., vertical Transformer blocks
II Related Work
In this section, we discuss related works, including HetNets, DRL, and transformer for wireless networks.
II-A Heterogeneous Networks
HetNets have been studied using a range of techniques, including optimization [17], game theory [18], heuristics [19], and more recently, DRL [20]. These approaches aim to address the diverse user demands and dynamic resource constraints inherent in HetNets. Despite progress, the high dimensionality and coupling among decision variables make HetNets a particularly challenging yet critical domain for intelligent network management [21]. In [22], the authors proposed a secrecy-aware, energy-efficient scheme for HetNets, jointly optimizing power control, channel allocation, and beamforming under secrecy and Signal-to-Interference-plus-Noise Ratio (SINR) constraints. They formulated the problem as a Markov Decision Process (MDP) and developed a multi-agent DRL framework, i.e., SecBoost, using dueling double DQN and prioritized experience replay to maximize long-term secrecy energy efficiency. In [23], a joint task offloading and resource management problem was formulated to improve computation efficiency in MEC-enabled HetNets. The authors addressed the resulting MINLP using an Advantage Actor-Critic (A2C)-based algorithm (A2C-JTRA), which jointly optimizes offloading decisions, transmit power, and CPU frequency for MEC servers and end devices. In [24], a secrecy-oriented joint optimization of beamforming, artificial noise, and power control was proposed for Tera Hertz (THz)-enabled femtocells. The authors used a Multi-Agent Reinforcement Learning (MARL) framework and introduced two DRL-based methods, i.e., Deep Deterministic Policy Gradient-based Secrecy Maximization (DDPG-SM) and Asynchronous-Advantage-Actor-Critic based Secrecy Maximization (A3C-SM), to operate in continuous action spaces and enhance long-term secrecy performance. While these methods demonstrate promising results, they often struggle to generalize across dynamic user behaviors and high-dimensional state-action spaces.
II-B DRL for Networks
DRL has been widely applied to network resource allocation tasks, where diverse user demands and time-varying conditions pose significant challenges to conventional optimization techniques [25]. In [26], the authors proposed a DRL-based Out-of-Order (OFO)-Friendly Multipath Scheduling (DOFMS) framework to ensure high-bandwidth and low-OFO transmission in mobile HetNets. The framework introduced a novel OFO metric for accurate evaluation, utilized a Double Deep Q-Network (DDQN) for dynamic scheduling, and incorporated an asynchronous learning module to handle variable path delays. In [27], a power control problem was formulated for spectrum-sharing HetNets to improve Global Energy Efficiency (GEE). The authors proposed a DRL-based cloud-edge collaborative framework that allows base stations to independently adjust transmit power, avoiding the scalability limitations of centralized optimization. In [28], a transformer-enhanced distributed DRL scheduling method was developed for heterogeneous IoT applications across edge and cloud computing environments. The method followed an actor-critic architecture, scaled across multiple distributed servers, and incorporated an off-policy correction mechanism to improve training stability. Despite these advances, existing DRL approaches often suffer from poor generalization and slow convergence in environments with heterogeneous user behaviors and high-dimensional state spaces.
II-C Transformer for Wireless Networks
Transformer-based models have emerged as a promising solution for complex wireless networks, offering strong reasoning capabilities and adaptability for high-dimensional decision-making tasks. In [29], the authors proposed a Decision Transformer (DT)-based architecture for wireless resource management, aiming to improve sample efficiency and generalization over conventional DRL. By pre-training in the cloud and fine-tuning at the edge, the DT framework achieved faster convergence and better performance in intelligent reflecting surface-aided and unmanned aerial vehicle-assisted MEC scenarios. In [30], the authors studied deep learning-based wireless interference recognition under non-cooperative electromagnetic attacks. They proposed a multi-modal feature extraction framework with a dual transformer module and adaptive gradient modulation to improve recognition accuracy, convergence speed, and computational efficiency. In [31], a transformer-based model was introduced for Channel State Information (CSI) prediction in high-mobility scenarios. Using multi-head attention and positional encoding, the model effectively captured global channel features and outperformed traditional deep learning models in accuracy with reasonable complexity. Transformers have also been used to approximate value functions in DRL, including Q-networks and critic networks. For example, the authors in [32] proposed the Deep Transformer Q-Network (DTQN), which employs transformer decoders to encode agent observation histories, yielding more stable and efficient Q-value estimation in partially observable environments. Similarly, the authors in [33] introduced a transformer-based Soft Actor-Critic (SAC) variant that uses transformers to process $N$-step returns for improved value estimation in complex tasks. The Transformer Off-Policy Episodic RL (TOP-ERL) method [34] similarly leverages transformers to estimate value functions from segmented action sequences, improving the handling of long-term dependencies. Other related works explore planning-inspired architectures. Value Iteration Networks (VIN) [35] embed differentiable planning procedures into neural networks, while the Predictron architecture [36] conducts internal rollouts to estimate multi-step returns. TreeQN and ATreeC [37] incorporate tree search mechanisms into value networks to enable forward-looking reasoning. While these approaches focus primarily on temporal modeling or multi-step return estimation, ReaCritic takes a different direction. It introduces a transformer-based critic specifically designed to enhance bidirectional reasoning within single-step value estimation. This design enables more expressive and accurate critic outputs in complex, dynamic environments without relying on full rollout-based return prediction.
III System Model
In this section, we present a HetNet system model with heterogeneous mobile users and formulate a multi-resource allocation problem, where users exhibit diverse preferences across different resource optimization objectives.
Figure 2: System model of the proposed HetNet with one BS and $M$ heterogeneous users characterized by distinct communication demands, computational capabilities, and service requirement.
III-A Wireless Environment
We consider a HetNet with $M$ heterogeneous mobile users, where each user represents either a real end user or a task-originating agent with specific communication demands, computational capabilities, and service requirements. As illustrated in Fig. 2, the network operates under dynamic and time-varying wireless conditions. The objective is to optimize joint resource allocation strategies, including transmission power, bandwidth, and computation resources, to adaptively satisfy user-specific QoS requirements.
To capture user mobility in a time-varying wireless environment under the coverage of the Base Station (BS), we adopt a random mobility model, where each user’s normalized position is updated at every time step via Gaussian perturbation [38]:
[…]
where $m\in\{1,\ldots,M\}$, $\Delta_{x_{m}}^{\left(t\right)}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ and $x_{m}^{\left(t\right)}$ is the normalized position of the $m_{\rm th}$ user at $t_{\rm th}$ step, and $\sigma$ controls the mobility intensity, allowing small random variations while ensuring that user positions remain within a bounded spatial domain. Each user’s transmission distance $D_{m}$ is derived from its normalized position and used to compute the large-scale path loss, modeled as
$\mathrm{\zeta}_{m}=D_{m}^{-\gamma},$ (2)
where $\gamma$ is the path loss exponent. For small-scale fading, we consider a Rician channel model, which accounts for both line-of-sight and multipath effects. Combining large-scale and small-scale fading, the uplink and downlink SINRs for the $m_{\rm th}$ user, i.e., $\mathrm{SINR}_{{\rm{u}},m}$ and $\mathrm{SINR}_{{\rm{d}},m}$ are defined as
$\mathrm{SINR}_{{\rm{u}},m}=\frac{p_{{\rm{u}},m}P_{{\rm{u}},{\rm{m}}}\cdot q_{{\rm{u}},m}\cdot\mathrm{\zeta}_{m}}{\sum_{j\neq m}p_{{\rm{u}},j}P_{{\rm{u}},\max}\cdot q_{{\rm{u}},j}\cdot\mathrm{\zeta}_{j}+\sigma_{\rm{u}}^{2}},$ (3)
and
$\mathrm{SINR}_{{\rm{d}},m}=\frac{p_{{\rm{d}},m}P_{{\rm{d}},\max}\cdot q_{{\rm{d}},m}\cdot\mathrm{\zeta}_{m}}{\sum_{j\neq m}p_{{\rm{d}},j}P_{{\rm{d}},\max}\cdot q_{{\rm{d}},j}\cdot\mathrm{\zeta}_{j}+\sigma_{\rm{d}}^{2}},$ (4)
where $p_{\mathrm{u},m}\in[0,1]$ and $p_{\mathrm{d},m}\in[0,1]$ denote the normalized power allocation factors for uplink and downlink transmission. $P_{\mathrm{u},\max}$ is the maximum uplink transmit power of the $m_{\rm th}$ user, while $P_{\mathrm{d},\max}$ is the maximum downlink transmit power determined by the BS. The terms $q_{\mathrm{u},m}$ and $q_{\mathrm{d},m}$ represent the Rician small-scale fading gains for uplink and downlink, and the corresponding noise powers are denoted by $\sigma_{\mathrm{u}}^{2}$ and $\sigma_{\mathrm{d}}^{2}$.
Based on the SINR expressions, we derive three key performance metrics for each users in the HetNet: uplink and downlink data rates, energy efficiency, and service latency.
III-A1 Data Rate
The achievable uplink and downlink data rates, i.e., $\mathcal{R}_{{\rm{u}},m}$ and $\mathcal{R}_{{\rm{d}},m}$, are computed using Shannon’s capacity formula:
$\mathcal{R}_{{\rm{u}},m}=b_{{\rm{u}},m}B_{{\rm{u}},{\rm{m}}}\log_{2}\left(1+\mathrm{SINR}_{{\rm{u}},m}\right),$ (5)
and
$\mathcal{R}_{{\rm{d}},m}=b_{{\rm{d}},m}B_{{\rm{d}},{\rm{m}}}\log_{2}\left(1+\mathrm{SINR}_{{\rm{d}},m}\right),$ (6)
where $b_{{\rm{u}},m},b_{{\rm{d}},m}\in[0,1]$ are the normalized bandwidth allocation factors, and $B_{\mathrm{u},m}$ and $B_{\mathrm{d},m}$ are the maximum available bandwidths for uplink and downlink transmissions.
III-A2 Energy Efficiency
The uplink and downlink energy efficiencies, denoted by $\eta_{\mathrm{u},m}$ and $\eta_{\mathrm{d},m}$ (in bits per Joule), are defined as the ratio of achievable data rate to the corresponding transmission power:
$\eta_{\mathrm{u},m}=\frac{\mathcal{R}_{\mathrm{u},m}}{p_{{\rm{u}},m}P_{{\rm{u}},\max}},$ (7)
and
$\eta_{\mathrm{d},m}=\frac{\mathcal{R}_{\mathrm{d},m}}{p_{{\rm{d}},m}P_{{\rm{d}},\max}},$ (8)
where the denominator represents the actual transmit power allocated to the $m_{\rm th}$ user for uplink and downlink, respectively.
III-A3 Service Latency
We define service latency $L_{m}$ as the total time required to transmit and process a task with computational demand $d_{m}$ (in CPU cycles). The first term captures transmission delay, while the second reflects task processing delay as
$L_{m}=\frac{{d}_{m}}{\mathcal{R}_{{\rm{u}},m}}+\frac{{d}_{m}}{c_{m}C_{\max}\cdot\rho_{m}},$ (9)
where $c_{m}\in[0,1]$ is the computational resource allocation ratio for the $m_{\rm th}$ user, $C_{\max}$ is the total computational capacity available in the HetNet, and $\rho_{m}$ denotes the user’s local computation efficiency, defined as
$\rho_{m}=\frac{1}{\kappa_{m}f_{m}},$ (10)
with $\kappa_{m}$ representing the effective switched capacitance coefficient determined by the device architecture, and $f_{m}$ denoting the local CPU frequency of the $m_{\rm th}$ user.
III-B Problem Formulation
The considered multi-resource allocation problem focuses on the HetNet scenarios where heterogeneous users have different preferences over different resource optimization objectives. We use different weight parameters to denote the different preferences. The $m_{\rm th}$ user has specific optimization preferences captured via a weighted utility function $Q_{m}$, reflecting priorities among uplink and downlink throughput, energy efficiency, and latency. The optimization objective is given as
$\max_{\left\{\bm{a}_{1},\ldots,\bm{a}_{M}\right\}}$ $\sum_{m=1}^{M}Q_{m}\left(\bm{a}_{m}\right),$ (11)
s.t. $0\leq p_{\mathrm{u},m}\leq 1,$
$0\leq p_{\mathrm{d},m}\leq 1,\quad\sum_{m=1}^{M}p_{\mathrm{d},m}\leq 1,$
$0\leq b_{\mathrm{u},m}\leq 1,\quad\sum_{m=1}^{M}b_{\mathrm{u},m}\leq 1,$
$0\leq b_{\mathrm{d},m}\leq 1,\quad\sum_{m=1}^{M}b_{\mathrm{d},m}\leq 1,$
$0\leq c_{m}\leq 1,\quad\sum_{m=1}^{M}c_{m}\leq 1,$
where
${\bm{a}}_{m}=\{p_{{\rm{u}},m},p_{{\rm{d}},m},b_{{\rm{u}},m},b_{{\rm{d}},m},c_{m}\}$ (12)
denotes the action vector for the $m_{\rm th}$ user, and $Q_{m}$ is
$Q_{m}=\;$ $\alpha_{m}^{(1)}\cdot\mathcal{R}_{{\rm{u}},m}\left(p_{{\rm{u}},m},\;b_{{\rm{u}},m}\right)+\alpha_{m}^{(2)}\cdot\mathcal{R}_{{\rm{d}},m}\left(p_{{\rm{d}},m},\;b_{{\rm{d}},m}\right)$
$+\alpha_{m}^{(3)}\cdot\eta_{{\rm{u}},m}\left(p_{{\rm{u}},m}\right)+\alpha_{m}^{(4)}\cdot\eta_{\rm{d},m}\left(p_{{\rm{d}},m}\right)$
$-\alpha_{m}^{(5)}\cdot L_{m}\left(p_{{\rm{u}},m},\;b_{{\rm{u}},m},\;c_{m}\right),$ (13)
where $\alpha_{m}^{(i)}$, $i\in{1,\ldots,5}$, are user-specific preference weights corresponding to uplink data rate, downlink data rate, uplink energy efficiency, downlink energy efficiency, and service latency, respectively. The constraints jointly ensure that each user’s power, bandwidth, and compute allocations remain within feasible per-user limits, while also satisfying global capacity constraints across all users.
IV Large Reasoning Transformer-based DRL Critic-model Scaling
In this section, we formulate the Markov Decision Process (MDP) for the DRL-based HetNet management and introduce ReaCritic, a large reasoning transformer-based critic-model scaling scheme designed to address key challenges in HetNets.
IV-A DRL Markov Decision Process Formulation
DRL offers a principled framework for sequential decision-making by enabling an agent to interact with an environment $\mathcal{E}$. At each discrete time step $t$, the agent observes a state ${\bm{s}}_{t}\in\mathcal{S}$ from the state space and selects an action ${\bm{a}}_{t}\in\mathcal{A}$ according to its learned policy. The environment responds with a reward signal $r_{t}^{\mathcal{E}}\in\mathbb{R}$, and transitions to a new state ${\bm{s}}_{t+1}$ based on a transition function $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. The learning objective is to optimize the expected cumulative reward by modeling the agent–environment interaction as a MDP.
  • Environment: The environment $\mathcal{E}$ captures the dynamics of HetNets, including user mobility, time-varying wireless channels, interference, and fluctuations in communication and computation resources. It enforces global constraints on power, bandwidth, and computational capacity. At each time step $t$, the DRL agent receives user-specific observations such as channel quality, task demand, latency requirement, position, computation capability, and user type. The environment computes uplink and downlink rates, evaluates task latency, and provides a multi-objective reward signal that reflects a weighted combination of throughput, energy efficiency, and service latency.
  • State: The state of the $m_{\rm th}$ user at time $t$ is represented by a high-dimensional vector:
    ${\bm{s}}_{t}^{m}=$$\{d_{m},L_{m},x_{m},f_{m},\tau_{m},P_{{\rm{u}},{\rm{m}}},q_{{\rm{u}},m},\mathrm{\zeta}_{m},\sigma_{\rm{u}},$
    $P_{{\rm{d}},{\max}},q_{{\rm{d}},m},\sigma_{\rm{d}},C_{\max},\rho_{m}\},$ (14)
    where $\tau_{m}\in\{0,1,\ldots,\tau_{\max}\}$ denotes the user type. Each user type corresponds to a distinct set of preferences over multiple objectives, encoded by weights $\alpha_{m}^{(i)}$ $\left(i=1,\ldots,5\right)$ in the utility function $Q_{m}$. These preference profiles are hidden from the network manager, introducing uncertainty into the learning process and increasing the complexity of policy adaptation.
    The global state at time $t$ is obtained by concatenating all user states as
    ${\bm{s}}_{t}=\left\{{\bm{s}}_{t}^{1},{\bm{s}}_{t}^{2},\ldots,{\bm{s}}_{t}^{M}\right\}$ (15)
    forming a joint high-dimensional input space that scales linearly with the number of users $M$.
    This high-dimensional representation captures a rich set of features that reflect both individual-level heterogeneity and global system dynamics, which are essential for making fine-grained, context-aware resource allocation and scheduling decisions in HetNets. However, such complexity also introduces significant challenges for DRL-based methods, including increased sample complexity, potential overfitting, and difficulties in generalization. Therefore, advanced policy and critic architectures are necessary to effectively model and extract useful patterns from this high-dimensional, structured state space.
  • Action The action vector for the $m_{\rm th}$ user is defined based on the key optimization variables in the objective formulation as (12).
    At each time step $t$, the DRL agent jointly selects the global action vector as
    $\bm{a}_{t}=\left\{\bm{a}_{t}^{1},\bm{a}_{t}^{2},\dots,\bm{a}_{t}^{M}\right\}.$ (16)
  • Reward The final reward for each user is composed of a weighted sum:
    $r_{m}$ $=\alpha_{m}^{(1)}\mathcal{R}_{u,m}+\alpha_{m}^{(2)}\mathcal{R}_{d,m}-\alpha_{m}^{(3)}b_{u,m}-\alpha_{m}^{(4)}b_{d,m}$ (17)
    $-\alpha_{m}^{(5)}p_{u,m}^{2}-\alpha_{m}^{(6)}p_{d,m}^{2}-\alpha_{m}^{(7)}\cdot\mathbb{I}\left(l_{m}>l_{\mathrm{th},m}\right),$
    where $\alpha_{m}^{(i)}$, $i\in\{1,2,3,4,5\}$ are task-specific weight coefficients. $l_{m}$ represents the total latency experienced by the $m_{\rm th}$ user, and $\mathbb{I}\left(l_{m}>l_{\mathrm{th},m}\right)$ is an indicator function penalizing violations of latency constraints. $l_{\mathrm{th},m}$ is the latency requirement for the $m_{\rm th}$ user.
This MDP formulation enables the DRL agent to learn preference-aware scheduling strategies in highly complex HetNet environments with heterogeneous users. The joint state and action spaces are high-dimensional due to per-user variability and system-wide coupling. For instance, with $M=50$ users and each user state vector comprising $14$ elements, e.g., task demand, position, fading, noise, and compute capacity, the global state dimension reaches $700$. Similarly, with five decision variables per user, the joint action space becomes $250$-dimensional. This scale, combined with non-stationary dynamics, latent user preferences, and inter-user interference, poses substantial challenges for policy learning, generalization, and sample efficiency.
IV-B ReaCritic Design
Figure 3: The architecture of the proposed ReaCritic, which integrates two-dimensional reasoning: horizontal expansion via HRea steps to enrich token diversity, and vertical abstraction via VRea steps to enhance hierarchical value representation.
ReaCritic introduces a two-dimensional reasoning architecture designed to enhance the expressiveness and generalization capability of the critic network. The first stage, denoted as Horizontal Reasoning (HRea), performs token-level expansion by generating multiple parallel state-action embeddings to simulate diverse trajectories or perspectives. This enriches the input space with latent behavioral variations and reduces sensitivity to local minima. The second stage, denoted as Vertical Reasoning (VRea), applies stacked Transformer blocks to process these tokens hierarchically, enabling deep abstraction and structured value propagation through multi-head self-attention and feed-forward layers.
As illustrated in Fig. 3, the HRea and VRea steps together enable ReaCritic to support fine-grained relational modeling and robust Q-value estimation in complex, high-dimensional HetNets. The whole procedure is provided in Algorithm 1.
Algorithm 1 ReaCritic-based DRL Framework
Input:
  • Environment: Markov environment $\mathcal{E}$, replay buffer $\mathcal{D}$
  • DRL Algorithm $\mathcal{A}$: discount factor $\gamma$, actor network $\pi_{\theta}$ (if have), target networks $Q_{\bar{\phi}}$, $\pi_{\bar{\theta}}$, learning rates $\eta_{Q}$, $\eta_{\pi}$, mini-batch size $M$
  • ReaCritic: critic network $Q_{\phi}$ with parameters $\phi$, HRea steps $H$, VRea steps $V$
1:Initialize policy network $\pi_{\theta}$, critic network $Q_{\phi}$, and replay buffer $\mathcal{D}$
2:for each training iteration do
3:     # Collect experience
4:     Reset environment and observe initial state $s_{0}$
5:     while not terminal do
6:         Sample action $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$ according to DRL algorithm $\cal A$’s policy or other algorithm-specific rules
7:         Execute $a_{t}$, observe $r_{t}$, $s_{t+1}$, and termination flag $d_{t}\in\{0,1\}$
8:         Store transition $(s_{t},a_{t},r_{t},s_{t+1},d_{t})$ in $\mathcal{D}$
9:         Update $s_{t}\leftarrow s_{t+1}$, $t\leftarrow t+1$
10:     end while
11:     # Update phase
12:     Sample mini-batch $\{(s_{t},a_{t},r_{t},s_{t+1})\}$ from $\mathcal{D}$
13:     (ReaCritic Update)
14:     ## Horizontal Reasoning:
15:     Compute state-action embedding $\mathbf{z}_{0}$ using (18)
16:     Expand into $H$ tokens using (19) and add exploration noise via (20)
17:     ## Vertical Reasoning:
18:     Refine tokens through $V$ transformer blocks using (21)
19:     Aggregate token representations via step-wise attention (22)
20:     Predict final Q-value using (23)
21:     (Critic loss and Update)
22:     Compute target Q-value:
$y_{t}=\begin{cases}r_{t}+\gamma Q_{\bar{\phi}}(s_{t+1},\pi_{\bar{\theta}}(s_{t+1})),&\text{actor-critic}\\ r_{t}+\gamma\max_{a^{\prime}}Q_{\bar{\phi}}(s_{t+1},a^{\prime}),&\text{Q-learning}\end{cases}$
23:     Compute critic loss: $\mathcal{L}_{Q}=\frac{1}{M}\sum_{i=1}^{M}(Q_{\phi}(s_{t},a_{t})-y_{t})^{2}$
24:     Update critic parameters: $\phi\leftarrow\phi-\eta_{Q}\nabla_{\phi}\mathcal{L}_{Q}$
25:     if actor $\pi_{\theta}$ exists then
26:         Compute policy gradient using (25)
27:         Update actor:  $\theta\leftarrow\theta+\eta_{\pi}\nabla_{\theta}J$
28:     end if
29:end for
Output: Trained DRL algorithm $\cal A$
IV-B1 Horizontal Reasoning Path
The core motivation behind the HRea design originates from the success of inference-time scaling in LLMs. In LLMs, performance improves when more intermediate tokens are generated during reasoning, e.g., Chain-of-Thought (CoT), enabling the model to engage in multi-step deduction before reaching a final output. While the critic in DRL is not tasked with next-token prediction, it performs complex mappings from high-dimensional state-action pairs to value estimates. These mappings can benefit from intermediate representations that act as latent reasoning anchors, similar in spirit to the token sequences in LLMs.
However, unlike LLMs that sequentially generate tokens, DRL critic networks operate under a single-shot evaluation paradigm. Therefore, to mimic multi-step reasoning, ReaCritic introduces the concept of horizontal reasoning expansion: instead of processing a single embedding, the model replicates the input state-action token multiple times and augments each copy with a unique positional embedding, effectively simulating parallel hypothetical reasoning paths.
Let the input state and action pair be $\bm{x}\in\mathbb{R}^{B\times d_{\bm{s}}}$ and $\bm{a}\in\mathbb{R}^{B\times d_{\bm{a}}}$ for batch size $B$. We denote $\bm{x}\|\bm{a}$ as the concatenation of the observation vector $\bm{x}\in\mathbb{R}^{B\times d_{\bm{s}}}$ and the action vector $\bm{a}\in\mathbb{R}^{B\times d_{\bm{a}}}$ along the feature dimension, resulting in a joint representation of size $\mathbb{R}^{B\times(d_{\bm{s}}+d_{\bm{a}})}$. This concatenated input captures both environment state and decision action, and serves as the input to the critic’s embedding module. We define the base embedding through a linear projection followed by normalization as
$\mathbf{z}_{0}=\mu\left(\mathbf{W}_{\text{embed}}\left[\bm{x}\|\bm{a}\right]\right)\in\mathbb{R}^{B\times d_{\rm h}},$ (18)
where $\mathbf{W}_{\text{embed}}\in\mathbb{R}^{(d_{\bm{s}}+d_{\bm{a}})\times d_{\rm h}}$ is the trainable projection matrix, and $\mu(\cdot)$ denotes layer normalization. To enable multi-path reasoning, this base embedding is then expanded into a sequence of $H$ horizontally parallel reasoning tokens as
$\mathbf{Z}^{(0)}=\left[\mathbf{z}_{0}+\mathbf{e}_{1},\ldots,\mathbf{z}_{0}+\mathbf{e}_{H}\right]\in\mathbb{R}^{B\times H\times d_{\rm h}},$ (19)
where $\mathbf{e}_{i}$ is the learnable positional encoding for the $i_{\text{th}}$ HRea step. To promote diversity and encourage exploration during training, an optional Gaussian noise term $\mathcal{N}(0,\sigma^{2})$ can be added to each token:
$\mathbf{Z}^{(0)}\leftarrow\mathbf{Z}^{(0)}+\mathcal{N}(0,\sigma^{2}).$ (20)
This horizontally extended token sequence is then passed to the subsequent vertical reasoning transformer layers, where latent value estimation is refined through attention-based abstraction. The key intuition behind HRea is to provide the critic network with multiple perspectives on the same decision point, enhancing robustness, exploration, and its ability to model high-order interactions in complex environments.
IV-B2 Vertical Reasoning Path
While HRea expands a single embedding point into multiple hypothetical reasoning paths, each token in this set still represents a shallow projection of the original state-action pair. To perform deeper reasoning and generate higher-order decision insights, we introduce VRea as a hierarchical processing stack.
This design is motivated by the idea of latent multi-step reasoning in LLMs [39], where deep transformer layers repeatedly refine representations in hidden space to support complex, context-dependent computation. Here, VRea does not explicitly produce symbolic CoT, but instead performs recursive abstraction within the vector space. This allows the critic to simulate structured inference trajectories by reprocessing the latent input across multiple transformer layers, enabling deeper exploration and more robust Q value modeling.
VRea also shares conceptual links with diffusion policy [40] and iterative policy refinement, where each step in the stack can be interpreted as an implicit planning operation conditioned on both prior layers and current observations. Compared with simple feed-forward architectures, the VRea stack increases reasoning depth without increasing parameter count proportionally, and facilitates richer representation learning through hierarchical feature composition.
Concretely, the output of HRea, denoted as $\mathbf{Z}^{(0)}\in\mathbb{R}^{B\times H\times d_{\rm h}}$, is processed by a stack of $V$ transformer blocks. Each transformer block $\Psi^{(v)}$ consists of a multi-head self-attention module followed by a feed-forward MLP layer, both wrapped with layer normalization and residual connections in a pre-norm configuration. This stacked vertical structure is implemented as
$\mathbf{Z}^{(v+1)}=\Psi^{(v)}\left(\mathbf{Z}^{(v)}\right),\quad v=0,\ldots,V-1,$ (21)
where $\Psi^{(v)}(\cdot)$ denotes the $v_{\rm th}$ transformer block. Each layer refines the horizontal token sequence by integrating long-range dependencies and capturing high-order interactions through attention and non-linear transformation. After $V$ layers, the final representation $\mathbf{Z}^{(V)}\in\mathbb{R}^{B\times H\times d_{\rm h}}$ is aggregated using attention over the $H$ horizontal reasoning steps as
$\hat{\mathbf{z}}=\sum_{i=1}^{H}\frac{\exp\left(\left(\mathbf{Z}^{(V)}\cdot\mathbf{w}_{a}\right)_{i}\right)}{\sum_{j=1}^{H}\exp\left(\left(\mathbf{Z}^{(V)}\cdot\mathbf{w}_{a}\right)_{j}\right)}\cdot\mathbf{Z}^{(V)}\left[:,i,:\right],$ (22)
where $\mathbf{w}_{a}$ is a learned projection. The final Q-value estimate is obtained as
$Q(\bm{x},\bm{a})=\mathbf{w}_{q}^{\top}\cdot\mu(\hat{\mathbf{z}}),$ (23)
where $\mathbf{w}_{q}\in\mathbb{R}^{d_{\rm h}}$ is a trainable projection vector used to map the final latent representation to a scalar Q-value.
This vertical reasoning mechanism enables ReaCritic to synthesize structured value estimates by recursively abstracting state-action pairs, resulting in improved generalization and temporal stability across complex HetNet scenarios.
IV-B3 Integration with DRL algorithms
Figure 4: Integration of ReaCritic into DRL algorithms. The part $1$ shows the reasoning-augmented SAC pipeline. Parts $2\sim 5$ illustrate compatibility with other actor-critic methods, including TD3, A3C, DDPG, and PPO.
As shown in Fig. 4, ReaCritic is designed as a plug-and-play critic module compatible with standard actor-critic DRL algorithms. While our implementation is based on the Soft Actor-Critic (SAC) framework [41], the design can be generalized to other value-based methods such as Twin Delayed Deep Deterministic (TD3) [42], Deep Deterministic Policy Gradient (DDPG) [43], and Asynchronous Advantage Actor-Critic (A3C) algorithms [44].
At each interaction step with the environment, the agent observes a state $\bm{s}_{t}$, samples an action $\bm{a}_{t}\sim\pi_{\theta}\left(\bm{s}_{t}\right)$, receives a scalar reward $r_{t}$, and transitions to a new state $\bm{s}_{t+1}$. The tuple $(\bm{s}_{t},\bm{a}_{t},r_{t},\bm{s}_{t+1})$ is stored in the replay buffer $\mathcal{D}$.
To estimate Q-values, ReaCritic first constructs a joint state-action embedding using (18), which is then expanded into $H$ horizontally reasoned tokens via (19) and (20). This horizontal sequence $\mathbf{Z}^{(0)}$ is refined through $V$ vertically stacked transformer blocks as described in (21). After vertical reasoning, the refined tokens are aggregated across reasoning steps using attention as in (22). The final Q-value is then produced by a lightweight output head as defined in (23).
The critic is trained by minimizing the Bellman residual [45] as
$\mathcal{L}_{Q}=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s}^{\prime})}\left[\left(Q_{\phi}(\bm{s},\bm{a})\!-\!\left(r\!+\!\gamma Q_{\phi^{\prime}}(\bm{s}^{\prime},\pi_{\theta}(\bm{s}^{\prime}))\right)\right)^{2}\right].$ (24)
The actor is updated to maximize the expected Q-value:
$\nabla_{\theta}J(\pi)=\mathbb{E}_{\bm{s}\sim\mathcal{D}}\left[\nabla_{\bm{a}}Q_{\phi}(\bm{s},\bm{a})\big{|}_{\bm{a}=\pi_{\theta}(\bm{s})}\cdot\nabla_{\theta}\pi_{\theta}(\bm{s})\right].$ (25)
By embedding inference-time reasoning into the critic, ReaCritic enhances learning stability, improves Q-value estimation accuracy, and accelerates policy convergence in complex wireless environments. The whole process is shown in Algorithm 1
IV-B4 Theoretical Complexity and Convergence Analysis
Introducing ReaCritic increases the expressive power of the critic network at the cost of additional computational complexity. Compared to standard feedforward critic models, which compute a single forward pass through an MLP, ReaCritic employs a Transformer-based dual-axis reasoning architecture with $H$ horizontal tokens and $V$ stacked reasoning layers.
The per-update computational complexity of ReaCritic is $\mathcal{O}(BH^{2}d_{\rm h}+BHVd_{\rm h}^{2})$, where $B$ is the batch size, $H$ is the number of HRea steps, $V$ is the number of vertical Transformer blocks, and $d_{\rm h}$ is the hidden dimension. The first term corresponds to the cost of multi-head self-attention across $H$ tokens per layer, while the second term captures the complexity of feed-forward transformations in each Transformer block. This structure scales linearly with $B$ and $V$, and quadratically with $H$ and $d_{\rm h}$. Compared to standard MLP-based critics with complexity $\mathcal{O}(Bd_{\rm h}^{2})$, the added cost remains moderate for lightweight settings with small $H$ and $V$, e.g., $H{=}4$, $V{=}3$, and can be efficiently batched on modern GPUs or edge accelerators. More importantly, ReaCritic is designed to support scaling along both the depth ($V$) and breadth ($H$) dimensions, enabling deployment of large reasoning-capable critics. This two-dimensional scaling path aligns with recent trends in inference-time compute augmentation and allows the model to serve both compact on-device learning and large-scale centralized training when compute resources permit.
Regarding convergence, although ReaCritic increases model depth and nonlinearity, its use of pre-normalized residual blocks and standard Transformer modules supports stable training. Since it is integrated into conventional DRL update rules, e.g., Bellman backup and policy gradients, the theoretical convergence properties of the base algorithms, such as SAC or TD3, remain valid, assuming the critic maintains universal approximation capacity. The added reasoning depth improves the expressiveness of the value function, which can reduce approximation bias and accelerate convergence. Additionally, stepwise attention and horizontal perturbation act as implicit regularization, improving stability and credit assignment in complex, non-stationary environments.
V Numerical Results
We evaluate the effectiveness and generalizability of ReaCritic in two settings: (i) a custom-designed HetNet environment simulating dynamic wireless conditions and heterogeneous user demands, and (ii) standard DRL control benchmarks from OpenAI Gym to assess cross-domain applicability. All experiments are conducted on a workstation running Ubuntu 20.04, equipped with an AMD EPYC 7543 32-core CPU and 4 NVIDIA RTX 4060 GPUs.
V-A Experimental Setup
V-A1 Wireless Benchmark Environment
We construct a high-dimensional wireless environment simulating a HetNet with $M\in\left[10,50\right]$ users. Each user is randomly placed within a normalized 2D area and exhibits independent mobility, modeled via Gaussian perturbations per time step. Channel conditions follow a Rician fading model combined with log-distance path loss, with transmission distances sampled from $\left[5,10\right]$ meters. Channel gains are updated at each episode to reflect temporal variation.
Each user is characterized by heterogeneous attributes, i.e., task demand, latency requirement, computation capability, channel quality, physical location, and a one-hot encoded type indicator. The maximum system capacities are set to $20$ Hz bandwidth, $10$ W transmission power, and $100$ computation units. At each decision step, the agent selects a five-dimensional action vector per user, controlling uplink/downlink power, bandwidth, and compute allocation. Actions are normalized and clipped to satisfy global resource constraints. The environment computes uplink and downlink rates, total task latency, and a multi-objective reward incorporating throughput, energy efficiency, resource usage, and latency satisfaction. This setting captures the key challenges of resource allocation in dynamic, multi-user HetNets.
V-A2 DRL Benchmark Environments
To evaluate the general applicability of ReaCritic, we conduct experiments across two categories of benchmark environments. First, we compare against standard DRL algorithms, i.e., SAC, DDPG, A3C, TD3, and PPO, within our custom wireless resource allocation environment. These serve as strong baselines for measuring learning stability, sample efficiency, and final policy performance. Second, to assess task generalizability beyond wireless systems, we integrate ReaCritic into these DRL algorithms and test them on three representative OpenAI Gym [46] control tasks with varying complexity:
  • MountainCarContinuous-v0: A low-complexity continuous control task where the agent must build momentum by swinging back and forth to drive a car up a steep hill.
  • HumanoidStandup-v4: A medium-complexity task where a quadruped robot learns to stand and maintain balance by coordinating multiple joints.
  • Ant-v4: A high-complexity locomotion task involving a multi-jointed humanoid robot that must learn stable standing and walking behaviors through high-dimensional action coordination.
These environments provide diverse and progressively challenging domains to validate ReaCritic’s ability to enhance value estimation and policy learning across different settings.
V-B Experiments Performance Analysis
V-B1 Performance Evaluation in HetNet with Varying User Density
We evaluate how ReaCritic-based SAC performs under different numbers of users in the proposed HetNet environment, and compare it with standard SAC in terms of convergence behavior and Q-value estimates.
(a) HRea step number is 1, VRea step number is 1.
(b) HRea step number is 4, VRea step number is 3.
(c) HRea step number is 8, VRea step number is 1.
(d) HRea step number is 8, VRea step number is 3.
(e) HRea step number is 12, VRea step number is 3.
(f) HRea step number is 20, VRea step number is 1.
(g) SAC.
(h) Impact of HRea and VRea step settings on final reward in ReaCritic-based SAC.
Figure 5: Comparison of training performance with block number to be $4$, and user number to be 10 across ReaCritic-based SAC method with varying numbers of HRea steps and VRea steps, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 10 users is also included.
(a) HRea step number is 1, VRea step number is 7.
(b) HRea step number is 4, VRea step number is 5
(c) HRea step number is 8, VRea step number is 1.
(d) HRea step number is 8, VRea step number is 3.
(e) HRea step number is 16, VRea step number is 3.
(f) HRea step number is 20, VRea step number is 1.
(g) SAC.
(h) Impact of HRea and VRea step settings on final reward in ReaCritic-based SAC with 20 users.
Figure 6: Comparison of training performance with block number to be $4$, and user number to be 20 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 20 users is also included.
(a) HRea step number is 4, VRea step number is 7.
(b) HRea step number is 8, VRea step number is 1.
(c) HRea step number is 8, VRea step number is 3.
(d) HRea step number is 12, VRea step number is 3.
(e) HRea step number is 16, VRea step number is 1.
(f) HRea step number is 20, VRea step number is 1.
(g) SAC.
(h) Impact of HRea and VRea step settings on final reward in ReaCritic-based SAC with 50 users.
Figure 7: Comparison of training performance with block number to be $4$, and user number to be 50 across ReaCritic-based SAC method, and standard SAC. The final reward ReaCritic-based SAC with different settings of HRea step and VRea step for 50 users is also included.
Fig. 5 compares the training performance of ReaCritic-based SAC under different HRea and VRea configurations with a small HetNet of $10$ users. Compared to the standard SAC baseline, ReaCritic consistently achieves higher and more stable cumulative rewards. When both HRea and VRea steps are small, e.g., $(4,3)$ or $(8,1)$, the model already outperforms standard SAC. Moreover, the critic output remains stable and well-aligned with the reward trajectory, indicating a coherent value function approximation. In contrast, standard SAC shows unstable critic behavior—initially rising but quickly collapsing—which weakens the learning signal and degrades policy performance. To analyze scalability, we evaluate medium-sized ($M=20$) and large-scale ($M=50$) HetNets in Fig. 6 and Fig. 7, respectively. As the number of users increases, the HetNet becomes more heterogeneous and the state-action space grows substantially, leading to a more complex learning environment. The results show that higher HRea and VRea settings, e.g., $(12,3)$, $(16,5)$, yield better final rewards in such settings. This validates our hypothesis that larger reasoning capacity, enabled by deeper horizontal and vertical reasoning, is needed to model the increased interaction complexity.
Across all user scales, ReaCritic-based SAC consistently outperforms standard SAC in both convergence speed and final performance. Furthermore, the Q-value curves in ReaCritic cases maintain stability throughout training, confirming that the structured inference paths constructed by the HRea and VRea mechanisms facilitate more reliable and expressive value estimation. These findings highlight ReaCritic’s scalability and robustness in DRL training for dynamic, multi-user HetNets.
V-B2 Generalization across DRL Algorithms and Benchmark Environments
To verify the versatility of ReaCritic, we integrate it with multiple DRL backbones (SAC, TD3, DDPG, PPO, A3C), and evaluate performance on three OpenAI Gym tasks with varying complexity: MountainCarContinuous-v0, HumanoidStandup-v4, and Ant-v4.
TABLE II: Final Episodic Return on HumanoidStandup-v4 With and Without ReaCritic
Algorithm With Without Gain
SAC 158.7k 152.9k +3.8%
DDPG 85.2k 31.5k +170.8%
A3C 57.8k 53.9k +7.3%
TD3 53.2k 42.1k +26.5%
PPO 52.9k 51.0k +3.7%
Table II presents the final episodic return of five representative DRL algorithms on the HumanoidStandup-v4 task, comparing variants with and without ReaCritic. Across all methods, integrating ReaCritic leads to consistent performance improvements, with gains ranging from modest (e.g., SAC and PPO: +3.8% and +3.7%) to substantial (e.g., DDPG: +170.8%). These results demonstrate ReaCritic’s strong generalizability across different DRL backbones and task complexities. In particular, actor-critic algorithms like DDPG and TD3 benefit significantly from ReaCritic’s reasoning mechanism, which stabilizes Q-value estimation and mitigates value degradation—a common challenge in environments with long-horizon planning and high-dimensional dynamics. Even methods like A3C and PPO, which are less reliant on precise value estimation, achieve better final returns when enhanced with ReaCritic. This highlights the critic’s role in improving temporal abstraction and training stability.
(a) MountainCarContinuous-v0 with SAC.
(b) HumanoidStandup-v4 with SAC.
(c) Ant-v4 with SAC.
(d) MountainCarContinuous-v0 with ReaCritic-based SAC.
(e) HumanoidStandup-v4 with ReaCritic-based SAC.
(f) Ant-v4 with ReaCritic-based SAC.
Figure 8: The comparison of the training performance with ReaCritic-based SAC method, and standard SAC under MountainCarContinuous-v0, HumanoidStandup-v4 and Ant-v4.
Fig. 8 shows that the ReaCritic-based SAC performs robustly across different environments, including standard GYM benchmarks. And it is observed that the more complex the environment is, the better ReaCritic-based SAC outperforms SAC. In the simple environment type such as MountainCarContinuous-v0, the performance of ReaCritic-based SAC and SAC are almost the same. In moderately complex environment such as HumanoidStandup-v4 and in high-complexity humanoid control environment Ant-v4, the ReaCritic-based SAC outperforms SAC and achieves higher rewards. And compare Fig. 8(b) and Fig. 8(e) with the results in Table. II, it shows that the proposed ReaCritic-based SAC algorithms achieve higher performance and more stable convergence than the other ReaCritic-based DRL algorithms and DRL benchmark algorithms.
V-B3 Ablation Study on Horizontal Noise Exploration in HRea
We evaluate the effect of horizontal token perturbation in HRea by removing the noise injection and comparing its impact on learning stability and final performance. One key design in ReaCritic is the addition of Gaussian noise to the horizontally expanded tokens to enhance generalization. To assess its utility, Fig. 9 presents the episode return of ReaCritic-based SAC with and without noise under a 50-user HetNet setting. Results show that noise injection consistently improves training stability and final reward, especially in large-scale networks where modeling diversity and reducing overfitting become more critical.
Figure 9: Reward performance for ReaCritic-based SAC with random noise and without noise when user number is $50$.
VI Conclusion
We proposed ReaCritic, a large reasoning transformer-based critic-model scaling framework that introduces reasoning capabilities inspired by LLMs into deep DRL for intelligent HetNet management. By combining HRea over diverse state-action sequences with VRea through stacked transformer blocks, ReaCritic overcomes the limitations of conventional shallow critics in complex, dynamic environments. Its modular design supports integration with various critic-based DRL algorithms, enabling flexible deployment across tasks. Extensive experiments demonstrate that ReaCritic enhances generalization, robustness, and convergence, offering a promising direction for intelligent and scalable network management.